# EHRAgent: Code Empowers Large Language Models for Complex Tabular Reasoning on Electronic Health Records

### v1

@@ D.2 Additional Error Analysis
$\diamond$ Context Length. This type of errors occur when the input queries or dialog histories are excessively long, exceeding the context length limit.
@@
Appendix E Additional Case Studies
We present additional case studies to showcase the effectiveness of each module in EHRAgent, including medical knowledge integration (Figure 7), code interface (Figures 8 and 9), and long-term memory (Figure 10).
Figure 7: Case study of medical knowledge injection in EHRAgent on MIMIC-III dataset. Given a question related to ‘aspirin ec’ and ‘venous cath nec’, EHRAgent effectively integrates knowledge about their potential location in the database and the identifiers required to associate them.
@@ Appendix E Additional Case Studies
Figure 10: Due to the constraints of limited context length, we are able to provide only a limited number of examples to guide EHRAgent in generating solution code. For a given question, the initial set of examples is pre-defined and fixed, which may not cover the specific reasoning logic or knowledge required to solve it. From the original examples on the left, none of the questions related to either ‘count the number’ scenarios or procedure knowledge. In contrast, when we retrieve examples from the long-term memory, the new set is exclusively related to ‘count the number’ questions, thus providing similar solution logic for reference.

### v3

@@ G.2 Additional Error Analysis
$\diamond$ Context Length. This type of error occurs when the input queries or dialog histories are excessively long, exceeding the context length limit.
@@
G.3 Additional Empirical Comparison of Primary Programming Languages
We conduct an additional analysis based on the empirical results (byond main results in Table 1) to further justify the selection of Python as our primary programming language.
Data Complexity.
The SPIDER (Yu et al., 2018) dataset, which is commonly used in SQL baselines (Pourreza and Rafiei, 2023), typically only involves referencing information from an average of 1.1 tables per question. In contrast, the EHRQA datasets we utilized require referencing information from an average of 1.9 tables per question. This significant gap in # tablesquestions indicates that EHRQA requires more advanced reasoning across multiple tables.
Sample Efficiency.
SQL-based methods require more demonstrations. As SQL occupies a relatively smaller proportion of training data, it is quite difficult for LLMs to generate valid SQL commands. Usually, the methods need at least tens of demonstrations to get the LLMs familiar with the data schema and SQL grammar. In EHRAgents, we only need four demonstrations as few-shot multi-tabular reasoning.
Environment Feedback.
DIN-SQL (Pourreza and Rafiei, 2023) establishes a set of rules to automatically self-correct the SQL commands generated. Nevertheless, these rules are rigid and may not cover all potential scenarios. While it does contribute to enhancing the validity of the generated SQL commands to some extent, DIN-SQL lacks tailored information to optimize the code based on different circumstances, resulting in a lower success rate compared to self-debugging and EHRAgent, which provide error messages and deeper insights.
Execution Time Efficiency.
We acknowledge that when handling large amounts of data, Python may experience efficiency issues compared to SQL commands. We have also observed similar challenges when working with the TREQS dataset, which contains a massive database with millions of records. However, in the MIMIC-III dataset, EHRAgent (avg. 52.63 seconds per question) still demonstrates higher efficiency compared to the state-of-the-art LLM4SQL method, DIN-SQL (Pourreza and Rafiei, 2023) (avg. 103.28 seconds per question). We will consider the efficiency of Python when dealing with large-scale databases as one of the important future directions.
Appendix H Additional Case Studies
We present additional case studies to showcase the effectiveness of each module in EHRAgent, including medical information integration (Section H.1), long-term memory (Section H.2), code interface (Section H.3), and rubber duck debugging module (Section H.4). In addition, we include Figures 13 and 14 to showcase the entire workflow of EHRAgent with accumulative domain knowledge and coding capability for EHR reasoning.
H.1 Case Study of Medical Information Integration
Figure 8 presents a case study of integrated medical knowledge based on a given query. When faced with a question related to ‘aspirin ec’ and ‘venous cath nec’, EHRAgent effectively summarizes and integrates relevant knowledge containing drug and procedure information, successfully reasoning and identifying the potential locations of the required information (tables and records) in the EHR database (e.g., ‘prescriptions’, ‘d_icd_procedures’), along with the detailed identifiers (e.g., ‘ICD9_CODE’, ‘HADM_ID’) needed to associate them.
Figure 8: Case study of medical information injection in EHRAgent on MIMIC-III dataset. Given a question related to ‘aspirin ec’ and ‘venous cath nec’, EHRAgent effectively integrates knowledge about their potential location in the database and the identifiers required to associate them.
H.2 Case Study of Long-Term Memory
Figure 9 presents a case study of updating few-shot demonstrations from long-term memory. Due to the constraints of limited context length, we are only able to provide a limited number of examples to guide EHRAgent in generating solution code. For a given question, the initial set of examples is pre-defined and fixed, which may not cover the specific reasoning logic or knowledge required to solve it. Using long-term memory, EHRAgent replaces original few-shot demonstrations with the most relevant successful cases from past experiences for effective plan refinement. For example, none of the original few-shot examples relate to either ‘count the number’ scenarios or procedure knowledge; after selecting from the long-term memory pool, we successfully retrieve more relevant examples, thus providing a similar solution logic for reference.
Figure 9: Case study of long-term memory in EHRAgent on MIMIC-III dataset. From the original few-shot examples on the left, none of the questions related to either ‘count the number’ scenarios or procedure knowledge. In contrast, when we retrieve examples from the long-term memory, the new set is exclusively related to ‘count the number’ questions, thus providing a similar solution logic for reference.
H.3 Case Study of Code Interface
Figures 10 and 11 present two case studies of harnessing LLMs as autonomous agents in a multi-turn conversation for code generation, in comparison to a natural language-based plan such as ReAct. From the case studies, we can observe that ReAct lacks a code interface, which prevents it from utilizing code structures for efficient action planning and tool usage. This limitation often results in a lengthy context for ReAct to execute, which eventually leads to a low completion rate.
@@ H.3 Case Study of Code Interface
H.4 Case Study of Rubber Duck Debugging
Figure 12 showcases a case study comparing the interactive coding process between AutoGen and EHRAgent for the same given query. When executed with error feedback, AutoGen directly sends back the original error messages, making slight modifications (e.g., changing the surface string of the arguments) without reasoning the root cause of the error. In contrast, EHRAgent can identify the underlying causes of the errors through interactive coding and debugging processes. It successfully discovers the underlying error causes (taking into account case sensitivity), facilitating accurate code refinement.
Figure 13: A complete version of case study in Figure 7 showcasing interactive coding with environment feedback.
Figure 14: Case study of the complete workflow in EHRAgent. With EHR metadata and tool definitions, EHRAgent (1) integrates medical information to locate the required tables/records, (2) retrieves relevant examples from long-term memory, (3) generates and executes code, (4) iteratively debugs with error messages until the final solution.